OpenStudy (anonymous):

How do you simplify this? logbase8(logbase3(logbase2^512))

OpenStudy (anonymous):

Maybe you can use the equation editor?

OpenStudy (anonymous):

start from the inside out what is \[\log_2(512)\] ? solve \[2^x=512\] first

OpenStudy (anonymous):

\[\log _{8}(\log _{3}(\log _{2}512))\]

OpenStudy (anonymous):

2^9

OpenStudy (anonymous):

Start from the way inside, like the math rule. Then work way outward.

OpenStudy (anonymous):

ok so now you are at \[\log _{8}(\log _{3}(\log _{2}512))=\log _{8}(\log _{3}(9))\]

OpenStudy (anonymous):

next solve \(3^x=9\) don't think too hard

OpenStudy (anonymous):

let me know when you get 2

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

ok so now comes the only hard part, (not that hard) computing \[\log_8(2)\]

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

bingo!

OpenStudy (anonymous):

since \[8^{\frac{1}{3}}=2\] you have \[\log_8(2)=\frac{1}{3}\] you are not doing that badly!

OpenStudy (anonymous):

My final answer would be that? & I'm sorry I just can't understand precal! :(

OpenStudy (anonymous):

lol you just did it what is not to understand? looks like you are psyching yourself out yes the answer is 1/3

OpenStudy (anonymous):

It's hard, i'm so use to having the calculator do the work which is why I don't know what to do without a calculator since we can't use the calculator for these problems :(

OpenStudy (anonymous):

But thank you so much! Once again :)

OpenStudy (anonymous):

yw